# Thermalization slowing down of weakly nonintegrable quantum spin dynamics

### v1

@@ Abstract
We study thermalization slowing down of a quantum many-body system upon approach to two distinct integrability limits. Motivated by previous studies of classical systems, we identify two thermalization time scales: one quantum Lyapunov time scale is extracted by quantifying operator growth in time in an appropriately defined basis, while another ergodization time scale is related to statistics of fluctuations of the time-evolved operator around its mean value based on the eigenstate thermalization hypothesis. Using a paradigmatic Quantum Ising chain we find that both timescales diverge upon approach to integrability. The relative strength of the divergence of the scales depends on the particular integrable limit. This allows us to define two different universality classes of quantum thermalization: short- and long-range networks.
Introduction— The study of integrable and chaotic dynamics in quantum systems is an area of active investigation with the goal of explaining the emergence of statistical mechanics in interacting quantum systems, among others. Multiple observables have been identified and studied as probes of integrable/chaotic dynamics. One class of such probes is based on spectral properties of the system [1, 2, 3, 4, 5, 6, 7, 8] rooted in the BGS and Berry-Tabor conjectures [1, 2] and the expected random matrix-like behavior of quantum systems [9, 10]. Another class of observables is based on the operator growth or state evolution under integrable/chaotic Hamiltonians. Observables such as Out-of-time-ordered-correlation functions (OTOCs)  [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], circuit complexity  [24, 25, 26, 27], operator size [28, 29, 30, 31] and Krylov complexity [32, 33, 34, 35, 36, 37, 38, 39] fall into this category.
@@
In quantum mechanical systems, the Eigenstate Thermalisation Hypothesis (ETH) [50, 51, 52, 53, 54, 55, 56] is often used to characterize thermalization. Within the purview of ETH, there are ergodization time-scales (e.g. the Thouless time)  [52, 57, 58] which have been explored in various systems. These time scales differ between chaotic and integrable systems, and their exact nature is studied extensively. Another active direction of investigation involves the notion of Lyapunov-like time scales for quantum systems. Operator growth serves as a potential path (via OTOCs) to define an appropriate spectrum [59, 60, 61]. Similarly, the spectral function is also used [62]. In quantum systems without well-defined classical limits (for example, spin-$1/2$ chains) the quantum Lyapunov spectra behave quite differently from classical spectra and may suffer from definition ambiguities [61, 60]. There are better-defined notions of the maximum Lyapunov exponent, which is typically extracted from the growth exponent of an appropriately defined observable. These include the exponent of the OTOC [63] and Krylov complexity [32, 37].
In this manuscript, we take the concept of time scales, originally developed for classical networks, to quantum mechanical systems near integrability. We introduce the notion of quantum networks near integrability and characterize them by studying the dynamics of conserved quantities of the limiting integrable Hamiltonian. We employ the operator growth approach (using Krylov complexity) to define an appropriate notion of Lyapunov time. Operator growth is captured through the lens of Krylov complexity, which describes the evolution of an operator on a minimal basis. We then use ETH principles to extract another time scale, which we coin the ergodization timescale (in analogy to classical systems). The system that we study is a prototypical 1D quantun Ising spin-$1/2$ chain. Near the integrable limits, these time scales are compared and used to identify two classes of quantum networks, which we call short- and long-range networks.
The model— The prototypical spin system we employ in order to characterize long and short-range networks is the Quantum Ising chain [64, 65] given by the following Hamiltonian
@@
where $J,g$ and $h$ are real numbers describing nearest-neighbour interaction, longitudinal and transverse magnetic field respectively. The $\sigma_{i}^{x,z}$ are Pauli matrices, describing spin-$1/2$ algebra. The system is in general non-integrable and has been extensively studied through various probes of quantum chaos [66, 67, 68, 69, 70, 71]. In the limits $g\rightarrow 0$ (TFIM) or $h\rightarrow 0$ (classical), the Hamiltonian (1) becomes integrable. We study the chaos and thermalization in the vicinity of these limits in order to observe and quantify the possible differences between the two limits.
@@
In classical systems near integrable limits, the short and long-range networks are defined with respect to the connectivity of the actions of the integrable system. The system is defined as a long-range network if the connectivity (defined through an appropriately defined coupling range) is extensive in the number of actions $N$. In a short-range network the connectivity of the actions (i.e. the coupling range) is independent of the number of actions.
@@
Inspired by the classical definition, we study the conserved quantities of the TFIM in the integrable limit. In the limit $h=0$, the spin chain becomes classical and the conserved quantities are local, with the simplest one being $\sigma^{z}_{i}$. The support of $\sigma^{z}_{i}$ does not scale with system size and neither is it connected with the other simple conserved quantities (i.e with $\sigma^{z}_{j\neq i}$). In the limit $g=0$, the spin chain is extensively connected while still being integrable. This is reflected by the fact that the conserved quantities are non-local [^1] Corresponding to the fermionic operators in the Jordan-Wigner representation of the system. Some of these operators correspond to simple symmetry operations. For example, the spin-flip operation $\sigma^{z}\rightarrow-\sigma^{z}$ for all spins leaves the Hamiltonian invariant. The corresponding conserved quantity is $\prod_{i=1}^{N}\sigma^{x}_{i}$. Other conserved quantities are similarly represented as extensive (non-local) combinations of the local spin matrices. The support of such quantities grows with system size.
@@
Thus, we define (by analogy to the classical definition [46, 47, 45]), the following two types of networks: (i) Quantum short-range network (SRN)—A system that is weakly nonintegrable, where the conserved quantities corresponding to the integrable part are connected by the nonintegrable perturbation in a short-ranged way, as observed from the operator dynamics defined by standard commutator relations. (ii) Quantum long-range network (LRN)—A system that is weakly nonintegrable, where the conserved quantities corresponding to the integrable part are connected in a non-local all-to-all way by the nonintegrable perturbation and their operator dynamics.
@@
In what follows, we probe the timescales and ergodicity properties of operators near the two respective limits and study time scales related to the ETH-like behavior of the observables. We then compare the divergence of these two time-scales – ETH-related and related to the operator growth – in the vicinity of the two limits.
@@
Krylov complexity— There exists a large class of observables that quantify operator growth under Hamiltonian dynamics. A common feature among most of these probes is the choice of a basis to expand the time-evolved operator. Once the basis is chosen, then appropriate expectation values are defined and evaluated, which then serve to distinguish between chaotic and integrable systems.
@@
One such probe is Krylov complexity [32]. The steps to evaluate Krylov complexity begin with generating a minimal basis  [^2] Minimal with respect to a cost function; which in this case is the Krylov complexity itself., which is called the Krylov basis. The Krylov complexity captures the average position of an operator in the minimal basis under the operator’s unitary evolution in time with the Hamiltonian $H$. The construction of the Krylov basis relies on an appropriately chosen norm in the Hilbert space of operators. We employ the infinite-temperature Hilbert-Schmidt norm for our analysis:
@@
known as Krylov complexity, grows exponentially with $t$ for chaotic dynamics: $K(t)\sim e^{2\alpha t}$. The exponent $\alpha$ captures the strength of the chaotic dynamics and is bounded from above by the Maldacena-Shenker-Stanford [63] bound. One can then define a natural time-scale for a chaotic system as $\alpha^{-1}$, which we denote as $T_{\lambda}$ throughout this manuscript, since it captures the growth of operators under the Hamiltonian $H$.
Ergodization time— Eigenstate thermalization hypothesis [50, 54] provides a powerful tool to study and characterize thermalization in quantum mechanical systems. It serves as a way to probe the chaotic or integrable dynamics of a system through evolution of operators and states. The essential statement of ETH can be encapsulated in the following equation
@@
We use the fluctuations of $R(t)$ to extract a timescale, that we refer to as the ergodization timescale $T_{E}$. The definition of this scale largely follows that in the classical case, discussed in Ref. 40. We choose a random initial state |$\ket{\psi}$ as the typical state and track the evolution of $\mathcal{O}(t)$. As the function $\langle\mathcal{O}(t)\rangle$ evolves it eventually starts to fluctuate around the mean value $\overline{O}$, going above and below the mean with time. This defines the passage times $t_{i}$ of the function across the mean value: $\langle\mathcal{O}(t_{i})\rangle=\overline{\mathcal{O}}$ [46, 47, 45, 44, 43, 42, 41, 40]. Time intervals between the two subsequent passages
@@
are called excursion times, since they reflect the time spent by the expectation value $\langle\mathcal{O}(t)\rangle$ away from its mean value. We distinguish additionally positive, $\tau_{i,+}$, and negative, $\tau_{i,-}$ excursions times for excursions above and below the mean value $\overline{\mathcal{O}}$. For a given system and operator, the collected excursion times $\tau_{i}$ obey some distribution. We use use the moments—mean and variance—of this distribution to extract an ergodization time-scale, as defined later and discussed in the Supplementary Material.
Results— We study numerically the dynamical properties close to integrability of the quantum short- and long-range networks respectively using the definitions and the methods outlined above. The Hamiltonian is given by Eq. (1) and we pick the two conserved quantities in the two integrable limits of the Hamiltonian
@@
$\mathcal{O}_{g\rightarrow 0}$ $=\prod_{i}^{N}\sigma^{x}_{i}$ (9)
as the time-evolved operators/observables whose dynamical properties we study. Other choices of operators yielded similar results. Naturally, at the integrable limits, the operators are conserved, their corresponding Lyapunov exponents are 00 and their Lyapunov times definhed through the Krylov complexity diverge. Similarly to the classical case, we are interested in quantifying the divergence of the Lyapunov and ergodization times (from the two probes) upon approaching the integrable limits.
@@
Figure 1: The Lyapunov time $T_{\lambda}$ for $N=8$ spins extracted from the linear growth of the Lanczos coefficients of time evolved oprtator $\mathcal{O}(t)$ in the Krylov basis as a function of $h$ for SRN and $g$ for LRN respectively. In both limits $T_{\lambda}$ shows a clear increase upon approach to the integrable limit.
@@
Figure 1 shows the Lyapunov times $T_{\lambda}$ extracted from the linear growth of the Lanczos coefficients in the Krylov basis of the operators (8-9) and plotted in the log-log scale as functions of the integrability breaking parameters $g$ (LRN) or $h$ (SRN). In both cases the observed behavior of $T_{\lambda}$ suggests a power-law divergence for small values of $g$ or $h$. The exponents are extracted for a system of $N=8$ spins via a linear fit of the data on the log-log scale
$T_{\lambda,SRN}=-0.13633*\ln(h)+0.39385$
$T_{\lambda,LRN}=-0.12832*\ln(g)-0.03919.$
@@
To study the thermalization properties, we collect the statistics of the excursion times $\tau_{\pm}$ (7) of the expectation value of the respective operators (8-9) in a random state (different choices of the state yielded similar results). We study the mean $\mu$ and variance $\sigma^{2}$ of the positive and negative excursion times (7). To define an appropriate ergodization time $T_{E}$, we study the relative behavior of mean $\mu$ and variance $\sigma^{2}$ by collecting $10^{4}$ excursions. Increasing this number does not change the moments significantly. The following behavior of the moments is observed, that follows rather closely the classical weakly nonintegrable systems: For SRN the values of $\sigma$ are $\approx 10^{4}$ times greater than the $\mu$ close to integrability. This suggests that the typical timescale of fluctuations is dominated by the distribution tail rather than its mean. A natural ergodization timescale is then defined as $T_{E}=\sigma^{2}/\mu$ [46, 47, 45, 44, 43, 42, 41, 40]. For LRN the values of the variance $\sigma$ are of the order of the mean $\mu$. This suggests that the dynamics receive comparable contributions from the mean as well as fluctuations of the crossing times. The ratio of variance and mean is approximately constant near the integrable limit and does not change considerably as $g$ is varied. We define ergodization time as the mean $T_{E}=\mu$, but the variance $\sigma$ is a valid choice as well.
@@
Figure 2: Comparison of the ergodization time $T_{E}$ and Lyapunov time $T_{\lambda}$ for the SRN (near-Ising limit) and LRN (near-Free limit) with fits. The linear curves have been shifted with appropriate constants.
@@
In Fig. 2 we compare the ergodization times $T_{E,\pm}$ for the SRN and LRN based on positive/negative excursion times with the Lyapunov time $T_{\lambda}$ obtained via the Krylov method. Similarly to the Laypunov time $T_{\lambda}$ the ergodization time $T_{E}$ also shows a power-law divergence with the decrease of the integrability breaking parameter. The linear fits of the data in the log-log scale are
$T_{E,LRN}(\tau_{+})=-0.72628*\ln(g)-0.26899$
$T_{E,LRN}(\tau_{-})=-0.74604*\ln(g)-0.303103$
$T_{E,SRN}(\tau_{+})=-2.34314*\ln(h)-0.23821$
$T_{E,SRN}(\tau_{-})=-1.95023*\ln(h)+0.19841$
@@
where $\tau_{\pm}$ stands for the timescales extracted from the positive and negative excursion times. Ergodization times extracted from positive and negative excursion times show similar scaling close to integrability.
A central observation is the different relative scaling of $T_{E}$ and $T_{\lambda}$ upon approach to integrability for SRN and LRN, which serves as a marker for short- and long-range networks, and is similar to that of the classical long- and short-range networks. In Fig. 2 $T_{E}$ becomes larger than $T_{\lambda}$ by several orders of magnitude for the short-range network (SRN) as integrability is approached. This is due to the fact that the decay exponent for $T_{E}$ is $\sim 20$ times the decay exponent for $T_{\lambda}$ in this case. Consequently as integrability is approached, the ergodization times diverges much faster than the Lyapunov times. A similar observation was also made for the classical SRN [45].
For the long-range network (LRN), $T_{E}$ remains comparable to $T_{\lambda}$ close to integrability. The decay exponent of $T_{E}$ is of the same order as the decay exponent for $T_{\lambda}$. The ergodization and Lyapunov times therefore diverge in a similar fashion upon approaching integrability.
This suggests a universality of thermalization slowing down in quantum many-body systems based on the emergence of two distinct network classes. The short-range networks, characterized by local conserved quantities in the integrable limit, show a much faster slowing down of thermalization as compared to operator growth. Long-range networks, characterized by non-local conserved quantities in the integrable limit, show a comparable slowing down of thermalization and operator growth.
Figure 3: $T_{E,\pm}/T_{\lambda}$ for LRN and SRN, plotted with respect to $T_{\lambda}/T^{*}_{\lambda}$.
This is further emphasized in Fig. 3, which shows the ratio $T_{E,\pm}/T_{\lambda}$ as a function of $T_{\lambda}/T^{*}_{\lambda}$. Here $T^{*}_{\lambda}$ is the maximum value of $T_{\lambda}$ for each (SRN, LRN) of the networks and is required for an effective comparison of the two networks, since the range of $T_{\lambda}$ observed in Fig. 1 is different for the SRN and LRN networks. It is evident from Fig. 3 that for the short-range network, $T_{E,\pm}/T_{\lambda}$ scales as $T_{\lambda}^{\alpha}$ with $\alpha\gg 1$, especially near $\frac{T_{\lambda}}{T^{*}_{\lambda}}=1$. For the long-range network, $T_{E,\pm}/T_{\lambda}$ scales approximately as $T_{\lambda}$.
@@
Conclusions— In this manuscript, we extended the universality classes of thermalization of classical weakly non-integrable systems to the case of weakly non-integrable quantum many-body systems. The classes are defined by the two timescales quantifying thermalization: one timescale comes from the Krylov complexity of operator growth. In the non-integrable regime, the K-complexity grows as $\exp(t/T_{\lambda})$, defining a Lyapunov timescale. Another timescale, the ergodization time $T_{E}$, is inspired by ETH and is defined through the statistics of the time intervals between consecutive crossings of the expectation of time-evolved operator $\langle\mathcal{O}(t)\rangle$ around its mean value $\overline{\mathcal{O}}$. $T_{E}$ is then defined through the appropriate moments of the intervals. For $\mathcal{O}$ conserved in the intergable limit, both timescales are expected to diverge.
@@
The short-range network (SRN) is defined by the locality of interaction between the conserved quantities (in the integrable limit) as integrability is weakly broken. We find that the distribution of the crossing intervals is fat-tailed in this case and therefore $T_{E}$ is defined as the ratio of the variance and the mean. Conversely, the long-range network (LRN) is defined by the non-locality of interaction between the conserved quantities upon weak integrability-breaking. Then the first two moments of the distribution of the crossing intervals are comparable and thus we define $T_{E}$ as the mean of the distribution.

### v2

@@ Abstract
We study thermalization slowing down of a quantum many-body spin system upon approach to two distinct integrability limits. Motivated by previous studies of classical systems, we identify two thermalization time scales: one quantum Lyapunov time scale is extracted by quantifying operator growth in time on an appropriately defined basis, while another ergodization time scale is related to the statistics of fluctuations of the time-evolved operator around its mean value based on the eigenstate thermalization hypothesis. Using a paradigmatic Quantum Ising chain we find that both timescales diverge upon approach to integrability. We investigate the relative strength of the divergence in the two limits and find that despite significant qualitative differences in the mechanism of integrability breaking, the timescales diverge in a similar fashion. This allows us to establish a universality of integrability breaking in quantum spin dynamics.
Introduction— The study of integrable and chaotic dynamics in quantum systems is an area of active investigation with the goal of explaining the emergence of statistical mechanics in interacting quantum systems, among others. Multiple observables have been identified and studied as probes of integrable/chaotic dynamics. One class of such probes is based on spectral properties of the system [1, 2, 3, 4, 5, 6, 7, 8] rooted in the Bohigas-Giannoni-Schmit (BGS) and Berry-Tabor conjectures [1, 2] and the expected random matrix-like behavior of quantum systems [9, 10]. Another class of observables is based on the operator growth or state evolution under integrable/chaotic Hamiltonians. Observables such as Out-of-time-ordered-correlation functions (OTOCs)  [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], circuit complexity  [24, 25, 26, 27], operator size [28, 29, 30, 31] and Krylov complexity [32, 33, 34, 35, 36, 37, 38, 39] fall into this category.
@@
In quantum mechanical systems, the Eigenstate Thermalisation Hypothesis (ETH) [50, 51, 52, 53, 54, 55, 56] is often used to characterize thermalization. Within the purview of ETH, there are ergodization time-scales (e.g. the Thouless time etc.) [52, 57, 58, 59, 60] which have been explored in various systems. These time scales differ between chaotic and integrable systems, and their exact nature is studied extensively. Another active direction of investigation involves the notion of Lyapunov-like time scales for quantum systems. Operator growth serves as a potential path (via OTOCs) to define an appropriate spectrum [61, 62, 63]. Similarly, the spectral function is also used [64]. In quantum systems without well-defined classical limits (for example, spin-$1/2$ chains) the quantum Lyapunov spectra behave quite differently from classical spectra and may suffer from definition ambiguities [63, 62]. There are better-defined notions of the maximum Lyapunov exponent, which is typically extracted from the growth exponent of an appropriately defined observable. These include the exponent of the OTOC [65] and Krylov complexity [32, 37].
In this manuscript, we extend the concept of time scales, originally developed for classical networks, to quantum mechanical systems near integrability. We introduce the notion of quantum networks near integrability and characterize them by studying the dynamics of conserved quantities of the limiting integrable Hamiltonian. We employ the operator growth approach (using Krylov complexity) to define an appropriate notion of Lyapunov time. Operator growth is captured through the lens of Krylov complexity, which describes the evolution of an operator on a minimal basis. We then use ETH principles to extract another time scale, which we coin the ergodization timescale (in analogy to classical systems). The system that we study is a prototypical 1D quantum Ising spin-$1/2$ chain. Near the integrable limits, the two time scales are compared. Their behavior is used to identify universal features of integrability breaking, by considering qualitatively different mechanisms of integrability breaking, which we call short and long-range networks by analogy with the classical case [40, 41, 42, 43, 44, 45, 46, 47].
The model— The prototypical spin system we employ in order to characterize long and short-range networks is the Quantum Ising chain (QIC) [66, 67] given by the following Hamiltonian
@@
where $J,g$ and $h$ are real numbers describing nearest-neighbour interaction, longitudinal and transverse magnetic field respectively. The $\sigma_{i}^{x,z}$ are Pauli matrices, describing spin-$1/2$ algebra. The system is in general non-integrable and has been extensively studied through various probes of quantum chaos [68, 69, 70, 71, 72, 73]. In the limits $g\rightarrow 0$ (Transverse Field Ising Model (TFIM)) or $h\rightarrow 0$ (Longitudinal Field Ising Model (LFIM)), the Hamiltonian (1) becomes integrable. We study chaos and thermalization time scales in the vicinity of these limits in order to observe and quantify the possible differences between the two limits.
@@
In classical systems near integrable limits, the way the actions are coupled by the integrability-breaking perturbations defines different classes of networks with different properties. The system is defined as a long-range network if the connectivity (defined through an appropriately defined coupling range) is extensive in the number of actions $N$. In a short-range network, the connectivity of the actions (i.e. the coupling range) is independent of the number of actions.
Inspired by the classical definition, we focus on the conserved quantities of the QIC in its integrable limits. In the limit $h=0$, the spin chain becomes effectively decoupled in real space and the conserved quantities are local, with the simplest one being $\sigma^{z}_{i}$. Adding a small nonzero value of $h$ will couple these conserved quantities in a local manner of a short-range network.
@@
In the limit $g=0$, the spin chain is extensively connected while still being integrable. This is reflected by the fact that the conserved quantities are non-local [^1] Corresponding to the fermionic operators in the Jordan-Wigner representation of the system. Some of these operators correspond to simple symmetry operations. For example, the spin-flip operation $\sigma^{z}\rightarrow-\sigma^{z}$ for all spins leaves the Hamiltonian invariant. The corresponding conserved quantity is $\prod_{i=1}^{N}\sigma^{x}_{i}$. Other conserved quantities are similarly represented as extensive (non-local) combinations of the local spin matrices or as sums of local terms. The support of such quantities grows with system size  [^2] Somewhat related to this is the delocalization in Fock space, which has been studied in Ref. 84.
@@
Thus, by analogy with the classical definition [46, 47, 45] we classify the quantum weakly non-integrable models by the character of coupling of conserved quantities in the integrable limit by the integrability breaking perturbation. We consider the following two types of networks: (i) Quantum short-range network (SRN)—a conserved quantity in the integrable limit is coupled by the integrability breaking perturbation to a system size independent number of other conserved quantities, as observed from the operator dynamics defined by standard commutator relations. (ii) Quantum long-range network (LRN)—a conserved quantity in the integrable limit is coupled by the integrability breaking perturbation to a number of other conserved quantities that scale with the system size.
@@
In what follows, we probe the ergodicity (ETH) and operator growth timescales of the above operators near the two respective limits. We then compare the divergence of these two time scales upon approaching each of the two limits.
Krylov complexity
@@
There exists a large class of observables that quantify operator growth under Hamiltonian dynamics. A common feature among most of these probes is the choice of a basis to expand the time-evolved operator. Once the basis is chosen, then appropriate expectation values are defined and evaluated, which then serve to distinguish between chaotic and integrable systems.
@@
One such probe is Krylov complexity [32]. The steps to evaluate Krylov complexity begin with generating a minimal basis  [^3] Minimal with respect to a cost function; which in this case is the Krylov complexity itself., which is called the Krylov basis. The Krylov complexity captures the average position of an operator in a minimal basis under the operator’s unitary evolution in time with the Hamiltonian $H$. The construction of the Krylov basis relies on an appropriately chosen norm in the Hilbert space of operators. We employ the infinite-temperature Hilbert-Schmidt norm for our analysis:
@@
known as Krylov complexity, grows exponentially with $t$ for chaotic dynamics: $K(t)\sim e^{2\alpha t}$. The exponent $\alpha$ captures the strength of the chaotic dynamics and is bounded from above by the Maladcena-Shenker-Stanford [65] bound. One can then define a natural time-scale for a chaotic system as $\alpha^{-1}$, which we denote as $T_{\lambda}$ throughout this manuscript since it captures the growth of operators under the Hamiltonian $H$.
Ergodization time— Eigenstate thermalization hypothesis [50, 54] provides a powerful tool to study and characterize thermalization in quantum mechanical systems. It serves as a way to probe the chaotic or integrable dynamics of a system through the evolution of operators and states. The essential statement of ETH can be encapsulated in the following equation
@@
are called excursion times, since they reflect the time spent by the expectation value $\langle\mathcal{O}(t)\rangle$ away from its mean value. We distinguish additionally positive, $\tau_{i,+}$, and negative, $\tau_{i,-}$ excursions times for excursions above and below the mean value $\overline{\mathcal{O}}$. For a given system and operator, the collected excursion times $\tau_{i}$ obey some distribution. We use the moments—mean and variance—of this distribution to extract an ergodization time scale, as defined later and discussed in the Supplementary Material.
Results— We study numerically the dynamical properties close to integrability of the quantum short- and long-range networks respectively using the definitions and the methods outlined above. The Hamiltonian is given by Eq. (1) and we choose the following conserved quantities in the two integrable limits of the Hamiltonian as the time-evolved operators/observables whose dynamical properties we study.
@@
$\mathcal{O}^{(k)}_{g\rightarrow 0}$ $=I^{(k)}\;\;\;\;k=1,\dots,N$ (9)
where $I^{(k)}$ denotes the set of conserved quantities for the Transverse field Ising model
$I^{(k)}$ $=iJ\sum_{j=1}^{N}\left(S^{zy}_{j:j+k}-S^{yz}_{j:j+k}\right)$ (10)
where we have the following shorthand
$S^{\alpha\beta}_{j:j+l}=\sigma^{\alpha}_{j}\left(\prod_{n=1}^{l-1}\sigma^{x}_{j+n}\right)\sigma^{\beta}_{j+l}$ (11)
The conserved quantity $\mathcal{O}^{(k)}_{h\rightarrow 0}$ is comprised of sums of $q$-local quantities, which have support on $q$ lattice sites. For $I^{(k)}$, we have $q=k+1$. At the respective integrable limits, the operators are conserved, their corresponding Lyapunov exponents are 00 and their Lyapunov times are defined through the Krylov complexity diverge. Similar to the classical case, we are interested in quantifying the divergence of the Lyapunov and ergodization times (from the two probes) upon approaching the integrable limits. For the following discussion, we will present the timescales obtained by averaging over the $N$ operators $\mathcal{O}^{(k)}_{g\rightarrow 0}$ in the LRN case. The individual timescales are presented in the Supplementary. For the SRN case, averaging over the $N$ possible $\sigma^{z}_{i}$ gives quantitatively similar timescales as that of an individual $\sigma^{z}_{i}$. This is discussed in the Supplementary.
@@
Figure 1: The Lyapunov time $T_{\lambda}$ for $N=8$ spins extracted from the linear growth of the Lanczos coefficients of time evolved operator $\mathcal{O}(t)$ in the Krylov basis as a function of $h$ for SRN and $g$ for LRN (averaged) respectively. In both limits $T_{\lambda}$ shows a clear increase upon approach to the integrable limit.
@@
Figure 1 shows the Lyapunov times $T_{\lambda}$ extracted from the linear growth of the Lanczos coefficients in the Krylov basis of the operators (8-9) and plotted in the log-log scale as functions of the integrability breaking parameters $g$ (LRN) or $h$ (SRN). Our LRN data for $T_{\lambda}$ are in semiquantitative agreement with similar computations in Ref. 32, although the Hamiltonian parameters do not completely match, and different operators were used.
In both LRN and SRN cases, the observed behavior of $T_{\lambda}$ is fitted with a (weak) power-law divergence for small values of $g$ or $h$. One may consider other fitting attempts for different model families in Ref. 32 which involve logarithmic fits. The differences are small in the considered parameter range, and all that matters for our purpose here is to use one and the same fitting procedure for all measured time scales. The exponents are extracted for a system of $N=8$ spins via a linear fit of the data on the log-log scale
$\log T_{\lambda,SRN}$ $=-0.083*\log(h)+0.218$ (12)
$\log T_{\lambda,LRN}$ $=-0.155*\log(g)+0.010.$ (13)
@@
To study the thermalization properties, we collect the statistics of the excursion times $\tau_{\pm}$ (7) of the expectation value of the respective operators (8-9) (denoted by the subscript $\sigma^{z}_{\text{SRN}}$ and $\langle I^{(k)}_{\text{LRN}}\rangle_{k}$ in the numerical plots) in a random state (different choices of the state yielded similar results). We average the timescales over the $N$ conserved quantities $I^{(k)}$ for the LRN case. The results for individual quantities $I^{(k)}$ (10) are discussed in the Supplementary. We study the mean $\mu$ and variance $\sigma^{2}$ of the positive and negative excursion times (7). To define an appropriate ergodization time $T_{E}$, we study the relative behavior of mean $\mu$ and variance $\sigma^{2}$, with the integrability breaking parameter, by collecting $10^{4}$ excursions. Increasing this number does not change the moments significantly. The following behaviour of the moments is observed, which follows rather closely the classical weakly nonintegrable systems: We observe that the values of $\sigma$ are exponentially larger than $\mu$ close to the integrable limit for both SRN and LRN, suggesting that the typical timescale of fluctuations is dominated by the distribution tail rather than its mean. A natural ergodization timescale is then defined as $T_{E}=\sigma^{2}/\mu$ [46, 47, 45, 44, 43, 42, 41, 40].
@@
Figure 2: Comparison of the ergodization time $T_{E}$ and Lyapunov time $T_{\lambda}$ for the SRN (near-Ising limit) and LRN (near-Free limit).
@@
In Fig. 2 we compare the ergodization times $T_{E,\pm}$ for the SRN and LRN based on positive/negative excursion times with the Lyapunov time $T_{\lambda}$ obtained via the Krylov method. Similarly to the Lyapunov time $T_{\lambda}$ the ergodization time $T_{E}$ also shows a power-law divergence with the decrease of the integrability breaking parameter. The linear fits of the data in the log-log scale are
$\log T_{E,SRN}$ $=-2.263*\log(h)-0.137$ (14)
$\log T_{E,LRN}$ $=-2.073*\log(g)-1.150$ (15)
@@
Ergodization times extracted from positive and negative excursion times $\tau_{\pm}$ show similar scaling close to integrability. Here we discuss the result obtained from $\tau_{+}$.
Our findings indicate that in the two integrable limits, the timescales $T_{\lambda}$ and $T_{E}$ diverge with exponents that differ by at least an order of magnitude. This follows from comparing (12) with (14) and (13) with (15). Therefore, the timescale associated with ETH diverges exponentially faster than that associated with operator growth. For the two network classes, the Lyapunov and Ergodization timescales respectively diverge with comparable exponents. This suggests a universality in the mechanism of integrability breaking in quantum many-body spin systems. The short-range network, characterized by local conserved quantities in the integrable limit demonstrates a slowing-down of thermalization at a rate which appears to be similar to the long-range network which is characterized by non-local conserved quantities (which appear as sums of local terms) in the integrable limit. However, such a conclusion can be partially deceiving, since we vary different parameters $g$ and $h$.
Figure 3: $T_{E}/T_{\lambda}$ for LRN and SRN, plotted with respect to $T_{\lambda}/T^{*}_{\lambda}$. The dashed lines represent the linear fit in the log-linear scale. The slopes correspond to $\alpha=6.16$ for LRN (in red) and $\alpha=12.63$ for SRN (in blue).
In order to properly compare the time scales from both network regimes, we replot them in units of the corresponding largest Lyapunov times. This is done in Fig. 3, which is our central result and shows the ratio $T_{E,\pm}/T_{\lambda}$ as a function of $T_{\lambda}/T^{*}_{\lambda}$. Now we put the time scale analysis of both classes on a similar footing  [^4] In particular, other forms of comparison have the underlying assumption that the integrability-breaking parameters $g$ and $h$ can be treated similarly. Comparison of $T_{E,\pm}/T_{\lambda}$ with $T_{\lambda}/T^{*}_{\lambda}$ frees us of this assumption.. Here $T^{*}_{\lambda}$ is the maximum value of $T_{\lambda}$ for each (SRN, LRN) of the networks and is required for effective comparison of the two networks, since the range of $T_{\lambda}$ observed in Fig. 1 is different for the SRN and LRN networks. It is evident from Fig. 3 that for the two network classes, $T_{E,\pm}/T_{\lambda}$ scales in qualitatively different ways as $T_{\lambda}^{\alpha}$ with $\alpha\gg 1$, especially near $T_{\lambda}/T^{*}_{\lambda}=1$. We find a rather weak scaling $\alpha=6.16$ for the LRN, as compared to a much stronger scaling $\alpha=12.63$ for the SRN.
@@
Conclusions— In this manuscript, we investigated the universality classes of thermalization of classical weakly non-integrable systems in the case of weakly non-integrable quantum many-body spin systems. The classes are defined by the two timescales quantifying thermalization: one timescale comes from the Krylov complexity of operator growth. In the non-integrable regime, the K-complexity grows as $\exp(t/T_{\lambda})$, defining a Lyapunov timescale.
Another timescale, the ergodization time $T_{E}$, is inspired by ETH and is defined through the statistics of the time intervals between consecutive crossings of the expectation of time-evolved operator $\langle\mathcal{O}(t)\rangle$ around its mean value $\overline{\mathcal{O}}$. $T_{E}$ is then defined through the appropriate moments of the intervals.
For $\mathcal{O}$ conserved in the integrable limit, both timescales are expected to diverge.
@@
The short-range network (SRN) is defined by the locality of interaction between the conserved quantities (in the integrable limit) as integrability is weakly broken. We find that the distribution of the crossing intervals is fat-tailed in this case and therefore $T_{E}$ is defined as the ratio of the variance and the mean.
Conversely, the long-range network (LRN) is defined by the non-locality of interaction between the conserved quantities upon weak integrability-breaking. Similar to the classical observation [45], we find that the two timescales respond in a qualitatively different way as compared to the SRN case, underscoring the difference of the integrability-breaking mechanisms.